integrate the following function. how to solve this type of questions
 1/(a+bsinx)

Grade:11

1 Answers

Aravind Suresh
7 years ago
First take 1/a common .=1/(1+b/a sinx)Now multiply and divide by (1-b/a sinx)Now the denominator is (b/a cosx)^2=» (1-b/a six)/(b/a cosx)^2One becomes secx^2 and other is tanxsecx.